OpenStudy (anonymous):

the mean height of 20 boys and 14 girls is 161cm.if the mean height of the 14 girls is 151cm,calculate the mean height of the 20 boys.

OpenStudy (anonymous):

Is this geometry @rids?

OpenStudy (anonymous):

no it is statistics

OpenStudy (anonymous):

Oh im sorry...I am in geometry..I dont know this....

OpenStudy (anonymous):

do you know any one who does?

OpenStudy (anonymous):

\[the mean height of 20 boys and 14 girls is 161cm.if the mean height of the 14 girls is 151cm,calculate the mean height of the 20 boys.\]

OpenStudy (anonymous):

let me check

OpenStudy (anonymous):

use what you know, how do you determine a mean?

OpenStudy (anonymous):

Well the simple way to do it,\[151+?\div2=161\]

OpenStudy (anonymous):

formula of mean=mean=sum of fx / f

OpenStudy (anonymous):

\[the mean height of 20 boys and 14 girls is 161cm.if the mean height of the 14 girls is 151cm,calculate the mean height of the 20 boys.\]

OpenStudy (anonymous):

The mean Heights of the Males would be 171 cm

OpenStudy (anonymous):

\[\large \frac{(b_1+b_2+...+b_{20})+(g_1+g_2+...+g_{14)}}{20+14}=N\] \[(b_1+b_2+...+b_{20})+(g_1+g_2+...+g_{14)}=34~N\] \[(b_1+b_2+...+b_{20})=34~N-(g_1+g_2+...+g_{14})\] \[\frac{(b_1+b_2+...+b_{20})}{20}=\frac{34}{20}~N-\frac{1}{20}(g_1+g_2+...+g_{14})\] \[\frac{(b_1+b_2+...+b_{20})}{20}=\frac{34}{20}~N-\frac{1}{20}~\frac{14}{14}(g_1+g_2+...+g_{14})\]

OpenStudy (anonymous):

\[boy_{\mu}=\frac{34(161)}{20}-\frac{14(151)}{20}\] looks about right to me

OpenStudy (anonymous):

I am in the eighth grade . do you have a simpler solution?

OpenStudy (anonymous):

math IS the simplest solution. and no, i cant think of a different way to approach that.

OpenStudy (anonymous):

thank u so much for all your help, I have another question.The heights of three plants A,B and C in a garden are in a ratio 2:3:5.their mean heights is 30cm.(a)find the height of plant B (b) if another plant D is added to the garden and the mean height of the four plants is now 33cm,find the height of plant d.

OpenStudy (anonymous):

let the heights be 2k, 3k, and 5k, and work the mean\[\frac{2k+3k+5k}{3}=30\] solve for k to know what 3k is equal to

OpenStudy (anonymous):

once you know k, we know the values for A,B C \[\frac{(A+B+C)+D}{4}=33\]solve for D since its the only unknown

OpenStudy (anonymous):

D = 33(4) - 10k, looks about right

OpenStudy (anonymous):

we are getting two variables D and K

OpenStudy (anonymous):

you there?

OpenStudy (anonymous):

k is from the setup in part a, once we know k we know the heights of A,B,C the setup in part b, uses the knowledge gained from part a, and all we have to solve for is D since k is known by then

OpenStudy (anonymous):

the mean of three numbers x,y and z is 6 and the mean of five numbers x,y,z,a,and b is 8. Find the mean of a and b.